# Latent Ewald summation for machine learning of long-range interactions

### v1

@@ III.3 Interfacial water
We employed the same settings for fitting the MLIPs as used for the bulk water dataset. The learning curves for forces are shown in Fig. 5. The energy errors are very low, reaching less than 0.1 meV/atom in mean absolute error (MAE) for all models using just 50 training configurations. The force errors are also small, ranging from about 14 meV/Å to 27 meV/Å in MAE when trained on 90% of the dataset. The learning behavior is similar to that observed in the bulk water case: both message passing and long-range interactions enhance the accuracy of the MLIPs.
Figure 6: A comparison of the interface properties of water predicted using short-range (SR) or long-range (LR) models and with or without massage passing layers ($T=0$ or $T=1$). a shows a snapshot of the thinner water slab configuration, b shows the water density profile, and c shows average cosine (c) the average $\cos(\theta)$ for the angle formed by the water dipole moment and z-axis. DFT results are from Ref. Niblett et al. (2021). d, e and f show the snapshot, water density, and $\langle\cos(\theta)\rangle$ for the thicker water slab, respectively.
@@
In addition to density profiles, we evaluated the orientational order profiles, shown in Fig. 6c,f, by computing the angle, $\theta$, between the dipole orientation of water molecules and the $z$-axis. For each model, the three different fits of MLIPs provide slightly different predictions, and the variances are larger for the SR models. At the interfaces, water molecules tend to form a dipole layer, which is screened by subsequent layers, ensuring that the bulk does not exhibit net dipole moments Niblett et al. (2021). However, without long-range interactions, this screening effect is not adequately captured, leading to extended dipole ordering into the bulk as seen in the $T=0$ SR model. The introduction of message passing ($T=1$ SR model) alleviates this issue at least for the thinner slab, but it also introduces an artifact: dipole ordering in the bulk along the opposite direction for the thicker slab. In contrast, the LR models recover the correct orientational order profiles, effectively capturing the screening effect and accurately representing the physical polarization behavior at the liquid-vapor interfaces.
IV Discussion and Conclusions
@@
In summary, we present a simple and general method, the Latent Ewald Summation (LES), for incorporating long-range interactions in atomistic systems. Unlike existing LR methods, LES does not require user-defined electrostatics or dispersion baseline corrections Niblett et al. (2021); Lee et al. (2022); Unke et al. (2021b), does not rely on partial charges or Wannier centers during training Unke and Meuwly (2019); Ko et al. (2021); Gao and Remsing (2022); Sifain et al. (2018); Gong et al. (2024); Shaidu et al. (2024), and does not utilize charge equilibration schemes Rappe and Goddard III (1991). We hypothesize that there may be underlying mathematical connections between the LES method and other approaches like LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) or recent density-based long-range electrostatic descriptors Faller et al. (2024), though LES offers a simpler mathematical formulation and implementation.
@@
Our results demonstrate that the LR model effectively reproduces correct dimer binding curves for pairs of charged, polar, and apolar molecules, long-range dipole correlations in liquid water, and dielectric screening at liquid-vapor interfaces. In contrast, standard short-ranged MLIPs, even when utilizing message passing, often yield unphysical predictions in these contexts. These challenges are common in atomistic simulations of molecules and materials, particularly in systems like water, which is ubiquitous in nature and essential to biological processes. Moreover, long-range effects are crucial in aqueous solutions, organic electrolytes, and other interfacial systems.
@@
We anticipate that the LES method will be widely adopted in atomistic modeling, especially for systems with significant electrostatic and dielectric properties. The long-range MLIP incurs only a modest computational cost, roughly double that of the short-range version (see Methods). Additionally, the LES method can be easily integrated into other MLIP frameworks, such as HDNNP Behler and Parrinello (2007), ACE Drautz (2019), GAP Bartók et al. (2010), MTPs Shapeev (2016), DeePMD Zhang et al. (2022), NequIP Batzner et al. (2022), MACE Batatia et al. (2022), etc.
V Methods
Implementation
@@ V Methods
We implemented the LES method as an Ewald module in CACE, which is written using PyTorch. The code is available in https://github.com/BingqingCheng/cace. The current implementation of the Ewald summation should in principle follow an overall scaling of $\mathcal{O}(N^{3/2})$ Toukmaji and Board Jr (1996). The implementation of the code can be further optimized by using algorithms with $\mathcal{O}(N\log(N))$ scaling Toukmaji and Board Jr (1996). The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent.
Benchmark of inference speed
@@
We benchmarked short-range and long-range CACE water models with and without message passing for MD simulations of liquid water using a single Nvidia L40S GPU with 48 GB of memory. Fig. 7 illustrates the time required per MD step. All models, with and without LR or message passing, exhibit near-linear scaling. SR models support simulations with up to approximately 30,000 atoms on a single GPU, while LR models handle around 10,000 atoms due to the higher memory demands of the Ewald routine, which may still be optimized. LR models run at roughly half the speed of SR models. It may be sufficient to use a lower $\mathbf{k}$-point cutoff in the LR model to further increase the speed.
@@
Figure 7: Timing of molecular dynamics simulations of the bulk liquid water system with varying number of atoms using different CACE models on a L40S GPU.
MD of Bulk Water
@@
To compute the longitudinal component of the dipole density correlation function for each model, we conducted an NVT simulation of 2,048 water molecules at 1 g/mL and 300 K in an elongated simulation box with a z-dimension of 99.3 Å. The simulation time for this system was 200 ps.
To compute the dipole density correlation functions, we calculated the dipole moment of each water molecule by assuming that the hydrogen atoms and oxygen atoms have charges of $+0.4238e$ and $-0.8476e$, respectively. This assumption only affects the absolute amplitude of the correlation function in Fig. 4, without altering the relative scale.
MD of Water Liquid-Vapor Interfaces
For the NVT simulations of the thinner slab at 300 K, we used a simulation cell with dimensions $(25.6~{}\mathrm{\AA},25.6~{}\mathrm{\AA},65.0~{}\mathrm{\AA})$, containing 522 water molecules. The thicker slab system had dimensions $(24.8~{}\mathrm{\AA},24.8~{}\mathrm{\AA},120.0~{}\mathrm{\AA})$, containing 1,024 water molecules. In both cases, the width of the water slab was less than half the total width of the cell. For the thinner slab, each independent NVT simulation lasted for about 400 ps. One run using $T=1$ SR model became unstable after about 280 ps and we discarded the unstable portion. For thicker slab, each run lasted for about 600 ps. One simulation using $T=1$ SR model became unstable after about 250 ps, and we only used the stable portion for the analysis.
@@
Competing Interests BC has an equity stake in AIMATX Inc.

### v2

@@ I.5 Example on interfacial water
We employed the same settings for fitting the MLIPs as used for the bulk water dataset. The learning curves for forces are shown in Fig. 5. The energy errors are very low, reaching less than 0.1 meV/atom in mean absolute error (MAE) for all models using only 50 training configurations. The force errors are also small, ranging from about 14 meV/Å to 27 meV/Å in MAE when trained on 90% of the dataset. The learning behavior is similar to that observed in the bulk water case: both message passing and long-range interactions enhance the accuracy of the MLIPs.
Figure 6: A comparison of the interface properties of water predicted using short-range (SR) or long-range (LR) models and with or without massage passing layers ($T=0$ or $T=1$). a shows a snapshot of the thinner water slab configuration, b shows the water density profile, and c shows average cosine (c) the average $\cos(\theta)$ for the angle formed by the water dipole moment and z-axis. DFT results are from Ref. Niblett et al. (2021). d, e and f show the snapshot, the water density, and $\langle\cos(\theta)\rangle$ for the thicker water slab, respectively.
@@
In addition to density profiles, we evaluated the orientational order profiles, shown in Fig. 6c,f, by computing the angle, $\theta$, between the dipole orientation of water molecules and the $z$-axis. For each model, the three different fits of MLIPs provide slightly different predictions, and the variances are larger for the SR models compared to the LR models. At the interfaces, water molecules tend to form a dipole layer, which is screened by subsequent layers, ensuring that the bulk does not exhibit net dipole moments Niblett et al. (2021). However, without long-range interactions, this screening effect is not adequately captured, leading to extended dipole ordering into the bulk as seen in the $T=0$ SR model. The introduction of message passing ($T=1$ SR model) alleviates this issue at least for the thinner slab, but it also introduces an artifact: dipole ordering in the bulk along the opposite direction for the thicker slab. In contrast, on average and with smaller variance between different fits , the LR models recover the correct orientational order profiles, effectively capturing the screening effect and accurately representing the physical polarization behavior at the liquid-vapor interfaces.
II Discussion
The current LES method can be extended in a number of ways. The form of the Ewald summation in Eqn. (4) was selected to best capture the electrostatic potential ($1/r$), which dominates long-range interactions. It remains unclear what is the effect of not directly accounting for other decaying potentials. Different forms can be used to capture other decaying potentials $1/r^{p}$ with different $p$ components. For examples, to best capture the London dispersion ($1/r^{6}$), one can use Chen et al. (1997); Williams (1971).
$E_{6}^{lr}=\left(\frac{\pi^{3/2}}{24V}\right)\sum_{k>0}k^{3}\\ \left[\pi^{1/2}\mathrm{erfc}(b)+\left(\frac{1}{2b^{3}}-\frac{1}{b}\right)\times\exp(-b^{2})\right]|S(\mathbf{k})|^{2},$ (5)
where $b^{2}=\sigma^{2}k^{2}/2$, and $\mathrm{erfc}$ denotes the complimentary error function. Currently, the hidden variable $q$ is rotational invariant, and it may be extended to adopt a vector or tensor form to capture the dipole and multiple moments on an atom.
The relation between LES and the LODE method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) is worth exploring. The key difference is the ordering of local and global interactions. In LES, the hidden variable $q_{i}$ is determined by the local atomic environment-dependent features $B_{i}$ of atom $i$, reflecting the nearsightedness principle: partial atomic charges depend mainly on local environments, similar to the short-range approximation in standard MLIPs. Although it should be pointed out that such locality approximation remains empirical rather than systematic. These local $q$ then interact globally via the Ewald summation to determine the global long-range energy. In contrast, in LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) the potential field (e.g. electrostatic potential and other $1/r^{p}$ decaying potentials) generated by all the atoms in the system is calculated in the reciprocal space via Ewald summation, and such field near a central atom $i$ up to some cutoff radius is then projected onto a set of basis functions to form the LODE descriptors. Finally, the short-range descriptors and the LODE descriptors of atom $i$ are used to predict its energy. The density-based long-range descriptors Faller et al. (2024) are similar to LODE, but the global atomic density itself is used instead of the field. In summary, LES uses local descriptors to predict global long-range energy, while LODE uses global field to inform local descriptors. There may be underlying mathematical connections between the LES and LODE, but this is not obvious, particularly considering the non-linearity introduced by the multilayer perceptron in LES that predicts $q_{i}$ from $B_{i}$ (Eqn. (2)).
LES also shares similarities with the recently proposed Ewald-based long-range message-passing method Kosmala et al. (2023), as both utilize structure factors. While LES directly determines the long-range energy from the structure factor of a hidden low-dimensional variable (Eqn.(4)), the Ewald message-passing method Kosmala et al. (2023) employs a learnable frequency filter to map the structure factor of atomic descriptors to real space. This mapping generates a long-range message for each atom that is used to update atomic descriptors during the message passing step. LES can thus be interpreted as a minimalist version of Ewald message-passing, with only one long-range message passing layer, compressed message dimension, fixed frequency filter, and simplified readout. Such minimalist design makes the model light-weight, and allows for easier physical interpretation.
@@ II Discussion
In summary, we present a simple and general method, the Latent Ewald Summation (LES), for incorporating long-range interactions in atomistic systems. Unlike many existing LR methods, LES does not require user-defined electrostatic or dispersion baseline corrections Niblett et al. (2021); Lee et al. (2022); Unke et al. (2021b), does not rely on partial charges or Wannier centers during training Unke and Meuwly (2019); Ko et al. (2021); Gao and Remsing (2022); Sifain et al. (2018); Gong et al. (2024); Shaidu et al. (2024), and does not utilize charge equilibration schemes Rappe and Goddard III (1991). Moreover, LES shows best accuracy compared to other long-range method and message passing neutral networks in the example on molten salt.
@@
Our results demonstrate that the LR model effectively reproduces correct dimer binding curves for pairs of charged and polar molecules, long-range dipole correlations in liquid water, and dielectric screening at liquid-vapor interfaces. In contrast, standard short-ranged MLIPs, even when utilizing message passing, often yield unphysical predictions in these contexts. These challenges are common in atomistic simulations of molecules and materials, particularly in systems like water, which is ubiquitous in nature and essential to biological processes. Moreover, long-range effects are crucial in aqueous solutions, organic electrolytes, and other interfacial systems.
@@
We anticipate that the LES method will be widely adopted in atomistic modeling, especially for systems with significant electrostatic and dielectric properties. The long-range MLIP incurs only a modest computational cost, roughly double that of the short-range version (see Methods). Furthermore, the LES method can be easily integrated into other MLIP frameworks, such as HDNNP Behler and Parrinello (2007), ACE Drautz (2019), GAP Bartók et al. (2010), MTPs Shapeev (2016), DeePMD Zhang et al. (2022), NequIP Batzner et al. (2022), MACE Batatia et al. (2022), etc.
III Methods
III.1 Implementation
@@
We implemented the LES method as an Ewald module in CACE, which is written using PyTorch. The code is available in https://github.com/BingqingCheng/cace. The current implementation of the Ewald summation should in principle follow an overall scaling of $\mathcal{O}(N^{3/2})$ Toukmaji and Board Jr (1996), with $N$ being the number of atoms. The implementation of the code can be further optimized using algorithms with $\mathcal{O}(N\log(N))$ scaling employing techniques with the charges interpolated to a 3D grid such as the Particle-Particle Particle-Mesh method  Toukmaji and Board Jr (1996). The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent.
III.2 Selection of the smearing factor $\sigma$
For a system of point charges, the Ewald summation expression in Eqn. (4) implies that the long-range energy corresponds to the electrostatic energy of Gaussian distributed charges with $\sigma$ being the standard deviation, while the short-range term needs to account for the difference between the Gaussian charges and the point charges. The value of the smearing factor $\sigma$ needs to be a fraction of the real-space cutoff radius $r_{\mathrm{rut}}$ in order to effectively eliminate the short-range error Kolafa and Perram (1992). Meanwhile, a larger $\sigma$ makes the Ewald series in $\mathbf{k}$ (Eqn. (4)) more converged and avoids a high reciprocal space cutoff $k_{c}$. and its optimal choice has been extensively discussed in the literature Kolafa and Perram (1992). Combining these considerations, $\sigma$ values in the rough range of 0.5 Å and 2 Å are reasonable choices.
We tested the choice of $\sigma$ on the bulk water dataset, with same settings as before and no message passing layer. Fig. 7 shows that $\sigma=1$ in this case leads to the lowest error, although the differences in outcome between various $\sigma$ choices are quite small.
Figure 7: Energy ($E$) and force ($F$) mean absolute errors (MAEs) on the bulk water dataset Cheng et al. (2019) with difference values of the smearing factor $\sigma$, using short-range models without massage passing.
III.3 Benchmark of inference speed
@@
We benchmarked short- and long-range CACE water models with and without message passing for MD simulations of liquid water using a single Nvidia L40S GPU with 48 GB of memory. Fig. 8 illustrates the time required per MD step. All models, with and without LR or message passing, exhibit favorable scaling. SR models support simulations with up to approximately 30,000 atoms on a single GPU, while LR models handle around 10,000 atoms due to the higher memory demands of the Ewald routine, which may still be optimized. LR models run at roughly half the speed of SR models. It may be sufficient to use a lower $\mathbf{k}$-point cutoff in the LR model to further increase the speed.
@@
Figure 8: Timing of molecular dynamics simulations of the bulk liquid water system with varying number of atoms ($N$) on a L40S GPU, using short-range (SR) or long-range (LR) CACE models and with or without massage passing layers ($T=0$ or $T=1$).
III.4 MD of Bulk Water
@@
To compute the longitudinal component of the dipole density correlation function for each model, we performed an NVT simulation of 2,048 water molecules at 1 g/mL and 300 K in an elongated simulation box with a $z$-dimension of 99.3 Å. The simulation time for this system was 200 ps.
To compute the dipole density correlation functions, we calculated the dipole moment of each water molecule assuming that hydrogen atoms and oxygen atoms have charges of $+0.4238e$ and $-0.8476e$, respectively. This assumption only affects the absolute amplitude of the correlation function in Fig. 4, without altering the relative scale.
III.5 MD of Water Liquid-Vapor Interfaces
For the NVT simulations of the thinner slab at 300 K, we used a simulation cell with dimensions $(25.6~{}\mathrm{\AA},25.6~{}\mathrm{\AA},65.0~{}\mathrm{\AA})$, containing 522 water molecules. The thicker slab system had dimensions $(24.8~{}\mathrm{\AA},24.8~{}\mathrm{\AA},120.0~{}\mathrm{\AA})$, containing 1,024 water molecules. In both cases, the width of the water slab was less than half the total width of the cell. For the thinner slab, each independent NVT simulation lasted for about 400 ps. One run using $T=1$ SR model became unstable after about 280 ps and we discarded the unstable portion. For the thicker slab, each run lasted for about 600 ps. One simulation using the $T=1$ SR model became unstable after about 250 ps, and we only used the stable portion for the analysis.
@@ III.5 MD of Water Liquid-Vapor Interfaces
Acknowledgements
BC thanks David Limmer for providing the water slab dataset, and Carolin Faller for the NaCl dataset.
Competing Interests The author declares no Competing Non-Financial Interests but the following Competing Financial Interests: BC has an equity stake in AIMATX Inc.
Author Contributions BC designed and performed the study, and wrote the paper.